How many different 7-digit phone numbers can you make using the integers 0 thru 9? Assume all numbers can be repeated and used in any position and that order matters.
Answer #1

There are 10 different digits. Each of 7 digit can be selected in 10 ways so using fundamental principle of counting number of different 7-digit phone numbers can you make using the integers 0 thru 9 are

10\cdot 10\cdot 10\cdot 10\cdot 10\cdot 10\cdot 10=10,000,000

Answer: 10,000,000